Circular No. 7803

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

JUPITER I (IO) E. Lellouch, Observatoire de Paris, Meudon; G. Paubert, Institut de

Radio-Astronomie Millimétrique; J. I. Moses, Lunar and Planetary Institute; and N. M. Schneider, Laboratory for Atmospheric and Space Physics, write: “We observed Jupiter I (Io) on Jan. 14–19 with the IRAM 30-m radiotelescope at Pico Veleta, Spain. We detected two rotational lines of NaCl at 234.252 and 143.237 GHz, providing the first observation of sodium chloride in Io’s atmosphere. By comparison with SO_{2} lines detected simultaneously at 234.187 and 143.057 GHz, we roughly estimate a NaCl/SO_{2} mixing ratio of ∼ 0.1 percent. Atmospheric NaCl may constitute a source of the long-observed sodium cloud around Io and of the recently discovered chlorine ions in the Io plasma torus.”

OPTICAL TRANSIENT IN SEXTANS A. Becker, Lucent Technologies (LT), Bell Laboratories; M. Lopez-

Caniego, Universidad Autonoma de Madrid; D. Norman, Cerro Tololo Interamerican Observatory (CTIO); and D. Wittman, LT, on behalf of the Deep Lens Survey team, report the detection, in five 600-s V -band frames taken on Jan. 11.32 UT and in five 600-s B-band frames taken on Jan. 12.22 with the CTIO 4-m reflector, of an unknown, hostless, point-source transient (V = 22.0, B = 22.4) located at α = 10^{h}49^{m}32^{s}.82, δ = −4^{o}31′45′′.4 (equinox 2000.0). This object was not present in five 600-s V -band and five 600-s B-band images taken on 2001 Mar. 27.0 to limiting magnitudes of ≈ 25.5. Additional photometry from Cerro Tololo and also by J. P. Halpern and M. Jackson with the MDM Observatory 1.3-m telescope on Kitt Peak (denoted by HJ): Jan. 13.251, R = 22.2; 13.334, R = 22.1 (HJ); 13.428, V = 22.0 (HJ); 14.220, R = 22.2; 14.237, V = 21.9; 14.249, B = 22.0.

SUPERNOVA 2002B IN ANONYMOUS GALAXY A. Clocchiatti, Pontificia Universidad Católica de Chile, on behalf

of the SuperMacho microlensing survey team (IAUC 7791), reports on his spectroscopic observations of SN 2002B obtained on Jan. 18.2 UT with the Las Campanas Observatory 6.5-m Baade telescope (+ dual imager/spectrograph LDSS2): “A fully reduced spectrum (400–800 nm; resolution 1.2 nm) shows that SN 2002B is a type-Ia supernova more than two weeks after maximum light. The spectrum resembles that of SN 1989B some 17 days after maximum. The redshift of the parent galaxy, measured from Hα and [N II] (658.3 nm), is 0.143.”

2002 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green